if a certain amount of food will last 50 days for 200 soldiers, how long will it last if 50 soldiers are added at the end of 20 days?
Answer by ankor@dixie-net.com(22740)

if a certain amount of food will last 50 days for 200 soldiers, how long will it last if 50 soldiers are added at the end of 20 days?
:
50 * 200 = 10000 food days, total available food
:
Let d = no. of days for this scenario
200d + 50(d-20) = 10000
200d + 50d - 1000 = 10000
250d = 10000 + 1000
250d = 11000
d = 44 days